# Towards a Deeper Understanding of Reasoning Capabilities in Large Language Models

### v1

@@ 4.2.3 TwoArmedBandit
The optimal strategy in the two-armed bandit task relies on a basic count of past rewards from both machines. We observe that incorporating more complex prompting led to performance degradation compared to the baseline. For instance, Llama3-8B’s score drops from $40.35\,[37.45-41.65]$ to $34.00\,[30.00-35.00]$ with Reflection + Planner. Similarly, DeepSeek-R1-14B score drops from $41\,[40.55,41.40]$ to $32.05\,[29.00,33.00]$ with Reflection + Planner. Llama-70B’s is the only exception, which achieves a higher maximum score of $48.00$ when Reflection + Planner is enabled compared to the maximum Baseline score $41.90$. Performance relies on rapid identification of the higher-paying bandit. The Base agent, exemplified by Llama3-8b, quickly exploits positive rewards, in contrast to other strategies that have lower performance by prolonged, oscillating exploration despite negative feedback (Example 1). Further inspection shows that Reflection, Oracle, and Planner discourage completely ignoring any single arm in favour of sustained exploration. Reflection points out to "avoid pulling the same machine too frequently without exploring the other option", while the Planner recommends alternating sequences like “Pull slot machine 2 → Pull slot machine 1 → Pull slot machine 2”. This analysis reveals two complementary failure modes. First, Reflection, Oracle and Planner increase the agent prompt extensively, diluting important information and decreasing performance in smaller models. This is consistent with the idea that extraneous context reduces signal-to-noise ratio and adds distraction [26, 11]. Second, advanced reasoning methods encourage the agent to keep testing both arms even when one arm is clearly superior. We often observe the model overthinks [2, 24, 5] and reasons itself away from exploiting a profitable streak, leading to slower convergence.
@@
Example 1: The Llama3-8b agent, with Reflection and Oracle, keeps exploring both bandits despite negative reward from slot machine 1. Player Observation Step 1: A new round begins. Player Observation Step 2: You pulled slot machine 1, you received reward -1. Player Observation Step 3: You pulled slot machine 2, you received reward 1. Player Observation Step 4: You pulled slot machine 1, you received reward -1. Player Observation Step 5: You pulled slot machine 2, you received reward 1. Player Observation Step 6: You pulled slot machine 1, you received reward -1.
4.2.4 Rock Paper Scissors
The highest gain appears for Llama3.3-70B: improving from $22.20\,[21.55,25.05]$ to $30.00\,[27.00-37.00]$ with Reflection + Planner. Llama3-8B improves from $16.50\,[16.40-19.05]$ to $26.00\,[12.00-36.00]$ with Reflection + Oracle. Mistral-Nemo-12B and DeepSeek-R1-14B do not see much improvement. Smaller models, especially Llama3-8B and mistral-nemo-12b, show occasional high-scoring runs but inconsistent overall play. Thus, only Llama3.3-70B benefits from prompting strategies here; other models see little reliable benefit.
@@ 4.2.4 Rock Paper Scissors
Qualitative inspections show that the larger model demonstrates better planning capabilities and adaptation to opponent patterns. For instance, Llama3.3-70B with Reflection + Planner reasoned that “choosing Paper has led to wins against Rock, which is a common choice”, while “being prepared to adapt with Scissor if the opponent chooses Paper, and then switch to Rock if the pattern suggests an upcoming Scissor.”. This ability to vary moves and execute planned sequences highlights a strong connection between its reasoning and actions. DeepSeek-R1-14B reflection shows a similar pattern, mentioning Track what has worked against the opponent and adjust your strategy dynamically, especially after losses or when the opponent changes their move. However, we note that DeepSeek-R1-14B often deviates from the answer template where the action line was missing, forcing random fallback. Smaller models often fell into repetitive loops. In one episode, Llama-8b agent opens with Paper and then continues to play Paper in most rounds, even while the opponent frequently plays Scissors. The agent incurs a string of losses but does not switch strategy. These cases show the smaller models failing to adjust even when a pattern clearly is not working, leading to predictable and exploitable behaviour.
4.2.5 Tower of Hanoi
@@
Llama3.3-70B achieves the best performance in Tower of Hanoi, with a score of 2.00 [2.00-2.00], matched by Reflection + Oracle. However, its score drops to 1.00 [1.00-2.00] with Reflection + Planner, and to 0.70 [1.00-1.80] with Reflection alone. We observe a trend in smaller models: while advanced prompting can significantly improve performance, this improvement is brittle, as it can also lead to substantial drops in scores. For instance, DeepSeek-R1-14B has a baseline score of 1.00 [0.60-1.20]. With Reflection + Planner, its score fluctuates, reaching as high as 2.00 but also dropping to 1.00. Similarly, for Llama3-8B, the baseline score of 0.20 [0.20-0.30] has lower variability, yet Reflection and Planner can achieve a maximum score of 2.00, though it may also drop to a minimum of 0.00. Mistral-Nemo-12B moves from 0.00[$0\,[0.00-0.00]0$ to $1.00\,[0.00-1.00]$ with both Reflection + Oracle and Reflection + Planner. However, the minimum score of $0.00$ is lower than that achieved with Reflection alone. Compared with the other games, agents using Planner struggle more to interpret the disk-placement state (see Supplementary Material for an ablation study on plan generation and accuracy). Hanoi failures frequently resulted from invalid moves, such as placing a larger disk onto a smaller one (Example 2). The agent’s reflection identified this recurring problem: "Based on the trajectory observed, there’s a common pitfall that has been repeatedly encountered: attempting to place a bigger disk on top of a smaller one.". While some models correctly state that Hanoi can be solved in 7 moves and can even list the sequence, our agents averaged around 30 moves without success, indicating significant inefficiency and a lack of true understanding and planning. Agents regularly made redundant actions such as moving a disk to a rod and then reversing that move or repeating failed actions. Adding to this difficulty was the sparse reward function: 0 for valid moves, -1 for invalid moves, and +100 for solving the puzzle, which provides little positive reinforcement while constantly being penalized for mistakes.
@@
Example 2: Mistral-Nemo-12b made an invalid move in Towers of Hanoi by attempting to place a larger disk (size 2 from rod B) onto a smaller one (size 0 on rod C). The game requires moving disks from A to C without placing a larger disk on a smaller one. State (Before Action): - A: |bottom, [], top| - B: |bottom, [2], top| - C: |bottom, [1, 0], top| Attempted Action: Move the top disk from rod B to rod C. Reward: -1 (Indicates an invalid move) Next State (Unchanged due to invalid action): - A: |bottom, [], top| - B: |bottom, [2], top| - C: |bottom, [1, 0], top|
4.2.6 Messenger
The biggest gain for smaller models comes from Reflection + Planner in Messenger. Llama3-8B’s score improves from 0.15[0.450.15)-0.15\,[-0.45--0.15)- 0.15 [ - 0.45 - - 0.15 ) to 0.00[$0\,[0.00-0.00]0$, Mistral-Nemo-12B improves from 0.20[0.200.15)-0.20\,[-0.20--0.15)- 0.20 [ - 0.20 - - 0.15 ) to 1.00[1.001.00]1.00delimited-[]1.001.001.00\,[-1.00-1.00]1.00 [ - 1.00 - 1.00 ], and DeepSeek-R1-14B improves from $0.40[0.25-0.55]$ to 1.00[0.401.00)1.00[0.40-1.00)1.00 [ 0.40 - 1.00 ). By contrast, the same Reflection + Planner collapses Llama3.3-70B from 0.10[0.100.10]0.10delimited-[]0.100.100.10\,[-0.10-0.10]0.10 [ - 0.10 - 0.10 ] to 1.00[1.001.00]1.00delimited-[]1.001.00-1.00\,[-1.00-1.00]- 1.00 [ - 1.00 - 1.00 ]. Qualitative analysis shows that this decline can be attributed to the planner’s strategy leading to overly cautious enemy avoidance, resulting in inefficient detours and failure to complete the task within ten steps. Object misidentification and poor spatial awareness frequently led to agent failures in Messenger. For example, Llama3-8b’s heuristics fail to adapt to the changing positions of enemy, message, and goal, often referring to objects that are not present and thereby confusing the agent. Mistral-Nemo-12b once incorrectly identified an airplane as the goal in one episode. Although the agent later attempted self-correction, stating, "the airplane, which is the enemy, is one step away and not moving. To avoid losing immediately, we should move towards the message (dog) located three steps southwest", its subsequent eastward move resulted in an enemy encounter. Even when the Oracle advises the agent "to navigate around the environment and avoid enemies while trying to reach the message", agents consistently struggled to move effectively. DeepSeek-R1-14B and Llama-70b demonstrate better synonym comprehension, and the latter is also better in creating generalizable heuristics: "Identify and move towards objects that match the description of a "message" (e.g., "classified report", "restricted document"), as obtaining the message is crucial to winning the game," and "Exercise caution around objects described with synonyms of "danger" or "enemy.". Similar to Hanoi, the reward function is not very informative. Positive outcomes only occur upon picking up the message or reaching the goal, events which happen infrequently. Most of the agent’s actions, such as wandering, offer no reward, and colliding with enemies ends the episode prematurely.
@@ 4.2.6 Messenger
Table 2: Mistral-Nemo-12b performance over five runs on Hanoi across four conditions and a random-action baseline. Values show the percentage of episodes where agents achieved the goal (G), average disks placed on target peg (D), and percentage of invalid moves (I). Darker colours mean better performance across the column. In the 3-disk task, showing valid actions proved more beneficial than reward shaping for disk placement and reducing invalid moves; combining them offered a minor boost to goal achievement. Surprisingly, the random policy outperforms all methods in the no adjustments setting across goal, disk placement, and invalid-move metrics. For 2 disks, all methods achieve substantially higher success than in the 3-disk scenario, with valid-action hints most effective at achieving the goal and reducing invalid moves. Combining shaping and showing valid actions yields no consistent advantage, indicating the main challenge is not the reward signal but identifying valid moves.
@@ 2 Disks
Table 3: Mistral-Nemo-12b performance over five runs on Messenger across four conditions and a random-action baseline. Values show the percentage of episodes where agents successfully picked up messages (P), achieved goals (G), or collided with enemies (C). Darker colours mean better performance across the column. Removing synonyms only slightly improves the pickup rates and the goal achievement. Reward shaping has a strong positive effect on the pickup rate but not on goal achievement. Combining both does not yield significant better results than reward shaping alone.
@@
4.3 Additional Analysis
@@
We modify the setups for Hanoi3Disks and MessengerL1 to investigate failure modes and explore potential learning enhancements in additional experiments. We conduct the experiments with Mistral-Nemo-12b to balance speed and performance.
4.3.1 Hanoi
@@
We implement the following modifications: simplifying the puzzle to two disks, mentioning the valid actions in the observation, and introducing reward shaping (-2 for invalid, +1 for valid moves, +100 for goal), to provide more explicit feedback. These modifications aim to isolate the factors hindering the agent’s capabilities. The results over five runs are summarized in Table 2.
@@
Under the baseline 3-disk setting, the agent never completes the puzzle and makes frequent illegal moves, typically ending episodes with only one disk on the target peg. Reward shaping provides some improvements. The agent places slightly more disks correctly and commits fewer illegal moves, but this does not lead to more goal achievements. Interestingly, showing valid actions reduces illegal moves but does not completely eliminate them, likely because these hints get lost in the lengthy prompt. The Planner strategy should map out correct steps, but often becomes confused by the puzzle’s intermediate states and stops choosing valid moves altogether. Combining reward shaping with valid-action hints yields no additional benefit beyond showing valid actions alone. Mistral-Nemo-12B underperforms even a uniform-random policy on success rate, correct placements, and move legality, revealing a fundamental ability to internalise the 3-disk Tower-of-Hanoi rules.
Under the two-disk scenario, all methods achieve much higher goal completion rates than in the three-disk setting. Without additional modifications, Reflection + Oracle attains 42 per cent goal success. Reward shaping generally boosts performance for most methods, with the exception of the Planner, while showing valid actions further reduces illegal moves and raises completion rates. However, combining reward shaping with showing valid actions does not consistently produce better outcomes like in the 3-disk setting, suggesting the main obstacle remains understanding the core constraint of the game.
4.3.2 Messenger
@@
Messenger modifications include reward shaping: providing incremental rewards for moving closer to the message or goal, and larger bonuses for message pickup (increased from 1.0 to 10.0) and final delivery (from 1.0 to 50.0). We introduce small distance-based rewards (+0.50 per step closer to the message or goal while carrying it), producing a denser reward signal to guide the agent. Second, we remove object synonyms to isolate the impact of language complexity on agent performance. The results over five runs are summarized in Table 3. Synonyms removal yields marginal gains in both goal completion and pickup rate for all methods, except for Reflection + Planner, where the pickup rate and goal achievement drop from 5.1 per cent to 4.9 per cent. Reward shaping consistently boosts pickup rates across all methods, but does not improve goal attainment, indicating that the agent either collides with the enemy or cannot attain the goal by the end of the episode. When reward shaping and synonym removal are combined, the pickup rate is higher and the collision is lower compared to the condition with no adjustments (except for the baseline). Reflection + Planner with reward shaping and synonym removal gives the best result overall (P = 47 percent, G = 8.5 percent, collisions = 27 percent). Despite reward shaping and synonym removal, high collision rates persist, indicating a fundamental limitation in spatial awareness and navigation.
5 Conclusion and Discussion

### v2

@@ IV-B3 TwoArmedBandit
The optimal strategy in the two-armed bandit task relies on a basic count of past rewards from both machines. We observe that incorporating more complex prompting led to performance degradation compared to the baseline. For instance, Llama3-8B’s score drops from 40.35[37.4541.65]40.35\,[37.45-41.65] to 34.00[30.0035.00]34.00\,[30.00-35.00] with Reflection + Planner. Similarly, DeepSeek-R1-14B score drops from 41[40.55,41.40]41\,[40.55,41.40] to 32.05[29.00,33.00]32.05\,[29.00,33.00] with Reflection + Planner. LLAMA3.3-70B is the only exception, which achieves a higher maximum score of $48.00$ when Reflection + Planner is enabled compared to the maximum Baseline score $41.90$. Performance relies on rapid identification of the higher-paying bandit. The Base agent, exemplified by Llama3-8b, quickly exploits positive rewards, in contrast to other strategies that have lower performance by prolonged, oscillating exploration despite negative feedback (Example 1). Further inspection shows that Reflection, Oracle, and Planner discourage completely ignoring any single arm in favour of sustained exploration. Reflection points out to ”avoid pulling the same machine too frequently without exploring the other option”, while the Planner recommends alternating sequences like “Pull slot machine 2 → Pull slot machine 1 → Pull slot machine 2”. This analysis reveals two complementary failure modes. First, Reflection, Oracle and Planner increase the agent’s prompt length extensively, diluting important information and decreasing performance in smaller models. This is consistent with the idea that extraneous context reduces signal-to-noise ratio and adds distraction [17, 18]. Second, advanced reasoning methods encourage the agent to keep testing both arms even when one arm is clearly superior. We often observe the model overthinks [19, 20, 21] and reasons itself away from exploiting a profitable streak, leading to slower convergence.
@@
Example 1: Llama3-8b with Reflection and Oracle keeps exploring both bandits despite negative reward from slot machine 1. Player Observation Step 1: A new round begins. Player Observation Step 2: You pulled slot machine 1, you received reward -1. Player Observation Step 3: You pulled slot machine 2, you received reward 1. Player Observation Step 4: You pulled slot machine 1, you received reward -1. Player Observation Step 5: You pulled slot machine 2, you received reward 1. Player Observation Step 6: You pulled slot machine 1, you received reward -1.
IV-B4 Rock Paper Scissors
@@
The highest gain appears for Llama3.3-70B: improving from 22.20[21.55,25.05]22.20\,[21.55,25.05] to 30.00[27.0037.00]30.00\,[27.00-37.00] with Reflection + Planner. Llama3-8B improves from 16.50[16.4019.05]16.50\,[16.40-19.05] to 26.00[12.0036.00]26.00\,[12.00-36.00] with Reflection + Oracle. Mistral-Nemo-12B and DeepSeek-R1-14B do not see much improvement. Smaller models, especially Llama3-8B and mistral-nemo-12b, show occasional high-scoring runs but inconsistent overall play. Inspections show that the larger model shows better planning skills and adaptation to opponent patterns. Llama3.3-70B with Reflection + Planner reasoned that “choosing Paper has led to wins against Rock, which is a common choice”, while “being prepared to adapt with Scissor if the opponent chooses Paper, and then switch to Rock if the pattern suggests an upcoming Scissor.”. This ability to vary moves and execute planned sequences highlights a strong connection between reasoning and action. Smaller models often fell into repetition. In one episode, Llama3-8B agent opens with Paper and then continues to play Paper in most rounds, even while the opponent frequently plays Scissors. The agent incurs a string of losses but does not switch strategy. These cases show the smaller models failing to adjust even when a pattern is not working, leading to exploitable behaviour.
IV-B5 Tower of Hanoi
@@
Llama3.3-70B achieves the best performance in Tower of Hanoi, with a score of 2.00 [2.00-2.00], matched by Reflection + Oracle. However, its score drops to 1.00 [1.00-2.00] with Reflection + Planner, and to 0.70 [1.00-1.80] with Reflection alone. While for smaller models, advanced prompting can significantly improve performance, this improvement is brittle and can lead to performance drops. DeepSeek-R1-14B has a baseline score of 1.00 [0.60-1.20]. With Reflection + Planner, its score fluctuates, reaching as high as 2.00 but also dropping to 1.00. Similarly, for Llama3-8B, the baseline score of 0.20 [0.20-0.30] has lower variability, yet Reflection and Planner can achieve a maximum score of 2.00, though it may also drop to a minimum of 0.00. Mistral-Nemo-12B moves from 0.00[0.000.00]0.00\,[0.00-0.00] to 1.00[0.001.00]1.00\,[0.00-1.00] with both Reflection + Oracle and Reflection + Planner. Compared with other tasks, agents using Planner struggle more to interpret the disk-placement state. Hanoi failures frequently resulted from invalid moves, such as placing a larger disk onto a smaller one (Example 2). The agent’s reflection identified this recurring problem: ”Based on the trajectory observed, there’s a common pitfall that has been repeatedly encountered: attempting to place a bigger disk on top of a smaller one.”. While the models can recite the optimal seven-move sequence for the Tower-of-Hanoi when prompted, the agents struggled, averaging 30 moves without completion, repeatedly shuffling disks back and forth or performing invalid moves. This illustrates a broader pattern: LLMs have seen Tower-of-Hanoi in training and can complete prompts about its optimal solution, yet they do not internalise the rules well enough to act on it. Another challenge is the sparse reward function: 0 for legal moves, -1 for illegal moves, and +100 upon completion, which gives the agent little positive feedback before the puzzle is solved.
@@
Example 2: Mistral-Nemo-12b made an invalid move in Tower of Hanoi by attempting to place a larger disk (size 2 from rod B) onto a smaller one (size 0 on rod C). State (Before Action): - A: —bottom, [], top— - B: —bottom, [2], top— - C: —bottom, [1, 0], top— Attempted Action: Move the top disk from rod B to rod C. Reward: -1 (Indicates an invalid move) Next State (Unchanged due to invalid action): - A: —bottom, [], top— - B: —bottom, [2], top— - C: —bottom, [1, 0], top—
@@
TABLE II: Mistral-Nemo-12b performance over five runs on Hanoi across four conditions and a random-action baseline. Values show the percentage of episodes where agents achieved the goal (G), average disks placed on target peg (D), and percentage of invalid moves (I). Darker colours mean better performance across the column. In the 3-disk task, showing valid actions proved more beneficial than reward shaping for disk placement and reducing invalid moves; combining them offered a minor boost to goal achievement. Surprisingly, the random policy outperforms all methods in the no adjustments setting across goal, disk placement, and invalid‑move metrics. For 2 disks, all methods achieve substantially higher success than in the 3-disk scenario, with valid‑action hints most effective at achieving the goal and reducing invalid moves. Combining shaping and showing valid actions yields no consistent advantage, indicating the main challenge is not the reward signal but identifying valid moves.
@@ 2 Disks
TABLE III: Mistral-Nemo-12b performance over five runs on Messenger across four conditions and a random-action baseline. Values show the percentage of episodes where agents successfully picked up messages (P), achieved goals (G), or collided with enemies (C). Darker colours mean better performance across the column. Removing synonyms only slightly improves the pickup rates and the goal achievement. Reward shaping has a strong positive effect on the pickup rate but not on goal achievement. Combining both does not yield significant better results than reward shaping alone.
@@
IV-B6 Messenger
Smaller models gain from Reflection + Planner in Messenger. Llama3-8B’s score improves from 0.15[0.450.15)-0.15\,[-0.45--0.15) to 0.00[0.000.00]0.00\,[0.00-0.00], Mistral-Nemo-12B improves from 0.20[0.200.15)-0.20\,[-0.20--0.15) to 1.00[1.001.00]1.00\,[-1.00-1.00], and DeepSeek-R1-14B improves from 0.40[0.250.55]0.40[0.25-0.55] to 1.00[0.401.00)1.00[0.40-1.00). The same Reflection + Planner collapses Llama3.3-70B from 0.10[0.100.10]0.10\,[-0.10-0.10] to 1.00[1.001.00]-1.00\,[-1.00-1.00]. This decline comes from the planner’s strategy, leading to overly cautious enemy avoidance, resulting in detours and failure to complete the task within ten steps. Object misidentification and poor spatial awareness frequently led to agent failures in Messenger. For example, Llama3-8b’s heuristics fail to adapt to the changing positions of enemy, message, and goal, often referring to objects that are not present and thereby confusing the agent. Mistral-Nemo-12b once incorrectly identified an airplane as the goal in one episode. Although the agent later attempted self-correction, stating, ”the airplane, which is the enemy, is one step away and not moving. To avoid losing immediately, we should move towards the message (dog) located three steps southwest”, its subsequent eastward move resulted in an enemy encounter. DeepSeek-R1-14B and LLAMA3.3-70B demonstrate better synonym comprehension, and the latter is also better in creating generalizable heuristics: ”Identify and move towards objects that match the description of a ”message” (e.g., ”classified report”, ”restricted document”), as obtaining the message is crucial to winning the game,” and ”Exercise caution around objects described with synonyms of ”danger” or ”enemy.”. Similar to Hanoi, the reward function is not very informative. Positive outcomes only occur upon picking up the message or reaching the goal, events which happen infrequently. Most actions offer no reward, and colliding with enemies ends the episode early.
IV-C Additional Analysis
@@
We modify the setups for Hanoi3Disks and MessengerL1 to investigate failure modes and potential improvements. The experiments are conducted with Mistral-Nemo-12b to balance speed and performance.
IV-C1 Hanoi
@@
We implement the following modifications: simplify the puzzle to two disks, add valid actions hints, and reward shaping (-2 for invalid, +1 for valid moves, +100 for goal). The results over five runs are summarized in Table II.
@@
Under the baseline 3-disk setting, the agent never completes the puzzle and makes frequent illegal moves, typically ending episodes with one disk on the target. With reward shaping, the agent places slightly more disks correctly and commits fewer illegal moves. Showing valid actions reduces illegal moves but does not eliminate them, likely because the hints get lost in the lengthy prompt. The Planner often becomes confused by the puzzle’s intermediate states and stops choosing valid moves altogether. Combining reward shaping with valid-action hints yields no benefit beyond showing valid actions alone. Mistral-Nemo-12B underperforms even a uniform-random policy on success rate, correct placements, and move legality, revealing a fundamental inability to internalise the 3-disk Tower-of-Hanoi rules. Under the two-disk scenario, all methods achieve higher goal completion rates than in the three-disk setting. Without additional modifications, Reflection + Oracle attains 42% goal success. Reward shaping boosts performance for most methods, except the Planner, while showing valid actions further reduces illegal moves and raises completion rates. However, combining reward shaping with showing valid actions does not consistently produce better outcomes, as in the 3-disk setting, indicating that LLMs can recall short move chains (2-disk) and that the actual bottleneck is the LLM’s lack of inherent understanding of the game’s core constraint rather than a shortage of feedback.
IV-C2 Messenger
@@
We apply reward shaping for Messenger: incremental rewards for moving closer to the message or goal, and larger bonuses for message pickup (increased from 1.0 to 10.0) and final delivery (from 1.0 to 50.0). We introduce small distance-based rewards (+0.50 per step closer to the message or goal while carrying it), producing a denser reward signal to guide the agent. We also remove object synonyms to isolate the impact of language complexity. The results over five runs are summarized in Table III. Synonyms removal yields marginal gains in goal completion and pickup rate for all methods, except for Reflection + Planner, where the pickup rate and goal achievement drop from 5.1% to 4.9%. Reward shaping consistently boosts pickup rates across all methods, but does not improve goal attainment. When reward shaping and synonym removal are combined, the pickup rate is higher, and the collision rate is lower than in the condition with no adjustments (except for the baseline). Reflection + Planner with reward shaping and synonym removal gives the best overall result (P = 47%, G = 8.5%, collisions = 27%). Despite reward shaping and synonym removal, high collision rates persist, indicating a fundamental spatial awareness and navigation limitation. Messenger exposes a gap between language comprehension and embodied competence: language-only pre-training equips LLMs with object labels and simple heuristics but not the grounded, stateful planning needed for dynamic spatial tasks. Bridging this gap will likely require tighter coupling between language and perception rather than prompt engineering.
V Conclusion and Discussion